# Anatomical Region Recognition and Real-time Bone Tracking Methods by Dynamically Decoding A-Mode Ultrasound Signals

### v1

@@ Abstract
Accurate bone tracking is crucial for kinematic analysis in orthopedic surgery and prosthetic robotics. Traditional methods (e.g., skin markers) are subject to soft tissue artifacts, and the bone pins used in surgery introduce the risk of additional trauma and infection. For electromyography (EMG), its inability to directly measure joint angles requires complex algorithms for kinematic estimation. To address these issues, A-mode ultrasound-based tracking has been proposed as a non-invasive and safe alternative. However, this approach suffers from limited accuracy in peak detection when processing received ultrasound signals. To build a precise and real-time bone tracking approach, this paper introduces a deep learning-based method for anatomical region recognition and bone tracking using A-mode ultrasound signals, specifically focused on the knee joint. The algorithm is capable of simultaneously performing bone tracking and identifying the anatomical region where the A-mode ultrasound transducer is placed. It contains the fully connection between all encoding and decoding layers of the cascaded U-Nets to focus only on the signal region that is most likely to have the bone peak, thus pinpointing the exact location of the peak and classifying the anatomical region of the signal. The experiment showed a 97% accuracy in the classification of the anatomical regions and a precision of around 0.5$\pm$1mm under dynamic tracking conditions for various anatomical areas surrounding the knee joint. In general, this approach shows great potential beyond the traditional method, in terms of the accuracy achieved and the recognition of the anatomical region where the ultrasound has been attached as an additional functionality.
@@ I INTRODUCTION
Bone tracking technology is essential for the kinematic analysis of the human body, particularly in the knee joint. The highly precise tracking produces accurate kinematics data, vital for surgical procedures [1], prosthetic robotics [2], and wearable exoskeletons [3]. Typically, the gold standard of tracking is achieved by using bone pins with optical markers [4], but it introduces invasive procedures and infection risks to subjects. Another method is electromyography (EMG)-based techniques [5, 6, 7, 8], but indirect measurement based on muscle activation patterns requires complex algorithms to analyze kinematics. In this context, a more accurate and convenient approach is preferable to obtain the knee kinematics in a non-invasive manner.
@@
Recently, an A-mode ultrasound (US) based tracking method has been introduced as a solution [9]. Compared to B-mode US, A-mode US can perform bone tracking in real-time, without the receiving and processing time of 2D images, and the need to analyze medical images by experts. Compared with other tracking techniques (e.g., bone pins, skin markers [10], fluoroscopy [11], MRI [12], etc.), A-mode ultrasound combines advantages of safety, high accuracy, non-invasiveness, and cost-effectiveness. However, its accuracy and robustness are compromised due to the reliance on traditional peak detection to analyze one-dimensional raw US signals. Traditional A-mode US methods ([4, 13, 9, 14]) used signal processing theory to detect the highest peak in the local range representing the precise bone distance from skin, which have not considered actual bone peak profiles, and heavily rely on the expert knowledge of the approximate local range. This paper tried to solve the issues with a deep learning framework.
In related research fields, deep learning has been employed for signal peak recognition. For example, in the diagnosis of heart disease, a U-Net framework was developed to segment and identify meaningful peaks in raw EEG signals [15]. However, the A-mode US signal presents a unique challenge: the meaningful bone peaks are actually sparse and ambiguous due to acoustic strength attenuation or the unclear interface between the soft tissues (e.g. tendon and muscles) and the bony surface. For this reason, to our knowledge, few studies have reported using deep learning for A-mode ultrasound diagnosis in knee kinematics tracking or real-time bone tracking. A novel method that considers both the local features (sparse bone peak) and the general features (entire ultrasound echo signal) of the A-mode US signal could be a potential solution to address this challenge.
@@
Figure 2: Our network structure had two U-Nets with different scales of perception fields for bone peaks detection and input signal classification. The input was a 1D ultrasound signal. The outputs were the signal classification results and the prediction of the peak location, which was a segment after thresholding the predicted probability sequence. The result Two U-Nets were connected by Sampling-based Proposal.
@@ II-A Motion Tracking System
In the end, the signals from different anatomical positions of knee joint were collected, with each anatomical region corresponded to one ultrasound channel, where the ultrasound transducer attached. All bone peak locations in the signals from all channels of femur and tibia epicondyle were annotated. These annotations served as dataset labels to train the proposed cascaded U-Nets separately (femur and tibia epicondyle). During testing, the network detected the bone peak in the signal, which was later converted to the actual depth of bone for evaluation using the calculated ground truth distance.
@@ II-E Evaluation
To evaluate our approach, we first collected the bias distance between the positions of predicted peaks (the peak position was regarded as the middle position of the segmentation) and the ground truth peaks. Then the mean and standard deviation of bias were calculated. Outliers that were much divergent from most biases were analyzed by examining the corresponding 3D position of the knee joint and the US waveform, which was shown in Fig. 3. We also recorded the network inference time to determine the speed of our method.
@@ III Result
TABLE I is the quantitative results of the bias distance. The dark background cells referred to the traditional method’s results. It showed that our method achieved an approximately $0.5\pm 1$mm the accuracy. In contrast, the traditional method using [14] showed the average accuracy of only $2.835\pm 1.893$mm. In addition, the cascaded UNets achieved high classification accuracies, with 97.04% for identifying three channels in femur epicondyle, and 97.48% for five channels in tibia epicondyle.
@@ IV Discussion
When looking at TABLE I, the bias between the predicted and ground truth peaks was around $0.5\pm 1$mm, indicating that the proposed method achieved a sub-millimeter accuracy for most cases, significantly surpassing the accuracy of the traditional method, as derived from [14]. In addition, there are slight variations across channels, possibly due to the different characteristics of the soft tissue surrounding the knee joint. However, these variations also provided unique signal characteristics beneficial for classification, resulting in a high accuracy rate of 97%. This highlights the efficacy of our anatomical-aware bone tracking approach, finely classifying anatomical regions during bone peak detection.
@@
In the future, this technique will be transferred and adopted in an in-vivo settings. In clinical practice, the bone-pins insertion for getting the ground truth labels is impossible due to the invasiveness. An alternative plan was to use other techniques (e.g., B-mode US [26], skin markers [10]) together to check the position of bony surfaces more precisely, and mark the approximate bone peaks in US signals to train the network. The further bone registration after using anatomical landmarks [9] can continually minimize the bone tracking errors. Additionally, since the bone peak profiles for in-vivo are different from the cadaver’s, new experiments using in-vivo’s data for training network are required to validate the accuracy of the approach, which is crucial to evaluate generalizability and robustness of method in clinical practice.

### v2

@@ Abstract
Accurate bone tracking is crucial for kinematic analysis in orthopedic surgery and prosthetic robotics. Traditional methods (e.g., skin markers) are subject to soft tissue artifacts, and the bone pins used in surgery introduce the risk of additional trauma and infection. For electromyography (EMG), its inability to directly measure joint angles requires complex algorithms for kinematic estimation. To address these issues, A-mode ultrasound-based tracking has been proposed as a non-invasive and safe alternative. However, this approach suffers from limited accuracy in peak detection when processing received ultrasound signals. To build a precise and real-time bone tracking approach, in this paper, a deep learning-based method was introduced for anatomical region recognition and bone tracking using A-mode ultrasound signals, specifically focused on the knee joint. The algorithm is capable of simultaneously performing bone tracking and identifying the anatomical regions where the A-mode ultrasound transducer was placed. It contains the fully connection between all encoding and decoding layers of the cascaded U-Nets to focus only on the signal region that is most likely to have the bone peak, thus pinpointing the exact location of the peak and classifying the anatomical region of the signal. The experiment showed a 97% accuracy in the classification of anatomical regions and a precision of around 0.5$\pm$1mm for tracking movements of the various anatomical areas surrounding the knee joint. In general, this approach shows great potential beyond the traditional method, in terms of the accuracy achieved and the recognition of the anatomical region where the ultrasound has been attached as an additional functionality.
@@ I INTRODUCTION
Bone tracking technology is essential for the kinematic analysis of human body, particularly in the knee joint. The highly precise tracking produces accurate kinematics data, vital for surgical procedures [1], prosthetic robotics [2], and wearable exoskeletons [3]. Typically, the gold standard of tracking is achieved by using bone pins with optical markers [4], but it introduces invasive procedures and infection risks to subjects. Another method is electromyography (EMG)-based techniques [5, 6, 7, 8], but indirect measurement based on muscle activation patterns requires complex algorithms to analyze kinematics. In this context, a more accurate and convenient approach is preferable to obtain the knee kinematics in a non-invasive manner.
@@
Recently, an A-mode ultrasound (US) based tracking method has been introduced as a solution [9]. Compared to B-mode US, A-mode US can perform bone tracking in real-time, without the receiving and processing time of 2D images, and the need to analyze medical images by experts. Compared with other tracking techniques (e.g., bone pins, skin markers [10], fluoroscopy [11], MRI [12], etc.), A-mode ultrasound combines advantages of safety, high accuracy, non-invasiveness, and cost-effectiveness. However, its accuracy and robustness are compromised due to the reliance on traditional peak detection to analyze one-dimensional raw US signals. Traditional A-mode US methods ([4, 13, 9, 14]) used signal processing theory to detect the highest peak (in a pre-defined local range) representing the precise bone distance from skin, which have not considered actual bone peak profiles, and heavily rely on the expert knowledge of approximate local range. This paper tried to solve the issues using a deep learning framework.
In related research fields, deep learning has been employed for signal peak recognition. For example, in the diagnosis of heart disease, a U-Net framework was developed to segment and identify meaningful peaks in raw EEG signals [15]. However, the A-mode US signal presents a unique challenge: the meaningful bone peaks are actually sparse and ambiguous due to acoustic strength attenuation or the unclear interface between the soft tissues (e.g. tendon and muscles) and the bony surface. For this reason, to our knowledge, few studies have reported using deep learning for A-mode ultrasound diagnosis in knee kinematics tracking or real-time bone tracking. A novel method that considers both the local features (sparse bone peak) and the general features (entire ultrasound echo signal) of A-mode US signals could be a potential solution to address this challenge.
@@
Figure 2: The proposed network had two U-Nets with different scales of perception fields (for bone peak detection) and input signal classification. The input was a 1D ultrasound signal. The outputs were the signal classification results and the prediction of the peak location. The peak location prediction was a segment after thresholding the predicted probability sequence. The two U-Nets were connected by Sampling-based Proposal.
@@ II-A Motion Tracking System
In the end, the signals from different anatomical positions of knee joint were collected, with each anatomical region corresponded to one ultrasound channel, where the ultrasound transducer attached. All bone peak locations in the signals from all channels of femur and tibia epicondyle were annotated. These annotations served as dataset labels to train the proposed cascaded U-Nets separately (femur epicondyle or tibia epicondyle). During testing, the network detected bone peaks in the signal, which was later converted to the actual depth of bone for evaluation using the calculated ground truth distance.
@@ II-E Evaluation
To evaluate our approach, we first collected the bias distance between the predicted peaks (the peak position was regarded as the middle position of the segmentation) and the ground truth peaks. Then the mean and standard deviation of bias were calculated. Outliers that were much divergent from most biases were analyzed by examining the corresponding 3D position of the knee joint and the US waveform, which was shown in Fig. 3. We also recorded the network inference time to determine the speed of our method.
@@ III Result
TABLE I is the quantitative results of the bias distance. The dark background cells referred to the traditional method’s results. It showed that our method achieved an approximately $0.5\pm 1$mm the accuracy. In contrast, the traditional method using [14] showed the average accuracy of only $2.835\pm 1.893$mm. In addition, the cascaded UNets achieved high classification accuracies, with 97.04% for classifying three channels signals in femur epicondyle, and 97.48% for classifying five channels signals in tibia epicondyle.
@@ IV Discussion
When looking at TABLE I, the bias between the predicted and ground truth peaks was around $0.5\pm 1$mm, indicating that the proposed method achieved a sub-millimeter accuracy for most cases, significantly surpassing the accuracy of the traditional method in the dark cells, derived from [14]. In addition, there are slight variations across channels, possibly due to the different characteristics of the soft tissue surrounding the knee joint. However, these variations also provided unique signal characteristics beneficial for classification, resulting in a high accuracy rate of 97%. This highlights the efficacy of our anatomical-aware bone tracking approach, finely classifying anatomical regions during bone peak detection.
@@
In the future, this technique will be transferred and adopted in an in-vivo settings. In clinical practice, the bone-pins insertion for getting the ground truth labels is impossible due to the invasiveness. An alternative plan was to use other techniques (e.g., B-mode US [26], skin markers [10]) together to check the position of bony surfaces more precisely, and mark the approximate bone peaks in US signals to train the network. The further bone registration after using anatomical landmarks [9] can continually minimize the bone tracking errors. Additionally, since the bone peak profiles for in-vivo are different from the cadaver’s, new experiments using in-vivo’s data for training network are required to validate the accuracy of the approach, which is crucial to evaluate generalizability and robustness of the method in clinical practice.